OpenStudy (anonymous):

Which of the following lists is ordered from least to greatest? −1.5, −2, 0, 2/5, 3/4 −2, −1.5, 0, 3/4, 2/5 −2, −1.5, 0, 2/5, 3/4 |−1.5|, −2, 0, 2/5, 3/4

OpenStudy (igreen):

B & C are the same

OpenStudy (mathstudent55):

What is |-1.5| = ?

OpenStudy (igreen):

A and D are wrong.

OpenStudy (mathstudent55):

Is there a number missing at the end of each line? Each line ends in two commas. Is there a square root, or some other symbol that did not get copied missing?

OpenStudy (anonymous):

fixed it

OpenStudy (mathstudent55):

Ok, that's better. I still have my question. What is |-1.5| = ?

OpenStudy (anonymous):

uh lol thats what i dont know..

OpenStudy (mathstudent55):

The symbol | | stands for absolute value.

OpenStudy (igreen):

Absolute value changes it to a positive number.

OpenStudy (igreen):

Which makes D, A WRONG ANSWER.

OpenStudy (mathstudent55):

Absolute value of a number is the number without a sign. |2| = 2 |0| = 0 |-3| = 3

OpenStudy (mathstudent55):

Now you can rewrite choice D with 1.5 instead of |-1.5|

OpenStudy (igreen):

For A: -1.5, 2 -1.5 > 2 So A is also WRONG.

OpenStudy (igreen):

B and C are about the same, but they have the fractions 2/5 and 3/4 in them. 2 / 5 = 0.40 3 / 4 = 0.75 So 2/5 < 3/4 Which means your answer is C.

OpenStudy (mathstudent55):

The next step is it would be helpful if you changed the fractions into decimals. What is 2/5 = ? What is 3/4 = ?

OpenStudy (mathstudent55):

Then rewrite each choice with the decimals instead of the fractions and the result of the absolute value. Then it will be easier to see if the numbers are in order.

OpenStudy (anonymous):

ok